# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v1

@@ 6 Experiments
We conduct experiments on five datasets of varying sizes and origins: Transaction financial dataset [agepred], MIMIC-IV medical dataset [johnson2020mimic], two social networks datasets, Retweet [zhao2015seismic] and StackOverflow [jure2014snap], and Amazon reviews dataset[jure2014amazon]. A detailed description of these datasets, along with the training and evaluation parameters, is provided in Appendix C.4.
@@
    Statistical baselines. MostPopular Generates a constant output with the most popular label from the prefix and the average time step. Last copies multiple previous events with adjusted timestamps.
@@
    Intensity-free approaches. We implement MAE-CE method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels [padhi2021tabgpt, mcdermott2024eventstreamgpt].
@@
    Intensity-based approaches. We implement RMTPP [du2016rmtpp] method as an example of the TPP approach with a traditional RNN. We also add NHP [mei2017nhp], based on a continuous time LSTM architecture.
@@
    Next-K approaches. We adapt MAE-CE and RMTPP for a direct prediction of multiple future events. This approach originates in time series analysis and has not been previously applied in the MTPP domain. The resulting methods are called MAE-CE-K and RMTPP-K, respectively.
@@
    Reranking. We also implement HYPRO [xue2022hypro], which generates multiple hypotheses and selects the best sequence using a contrastive approach.
Details on the methods can be found in Appendix A and Appendix C.1. Training specifics are provided in AppendixD. Hyperparameters for both methods and metrics are presented in Appendix E. The main evaluation results are presented in Table 1. Below, we will discuss these results from various perspectives and provide an additional analysis of the methods’ behavior.
6.1 OTD vs T-mAP
Evaluation results indicate that the OTD metric can yield low values even for simple statistical baselines. For instance, the Last baseline achieves the lowest OTD value on the Retweet dataset and outperforms Next-K methods on the Amazon dataset. In contrast, the T-mAP metric clearly differentiates between statistical baselines and deep models, as the former is incapable of predicting label probabilities or confidences.
@@ 6.2 Intensity-Based vs Intensity-Free
Previous works have compared either intensity-based neural methods [du2016rmtpp, xue2022hypro] or intensity-free approaches [padhi2021tabgpt, mcdermott2024eventstreamgpt]. This raises the question: which approach is superior? We compare the intensity-free MAE-CE method with intensity-based RMTPP and NHP approaches. Our results show that the intensity-free MAE-CE method excels in the next-event MAE prediction, as it optimizes this metric. In other scenarios, the results vary between datasets. Therefore, we conclude that both approaches warrant attention in future research.
@@ 6.3 Autoregression vs Direct Horizon Prediction
We compared MAE-CE and RMTPP methods with their Next-K counterparts. According to our results, autoregression generally performs better on Transactions, MIMIC-IV, and Amazon datasets. However, on social network datasets (Retweet and StackOverflow), Next-K models typically outperform autoregression in horizon prediction tasks. Despite this, little effort has been made to adapt Next-K models to MTPPs. Consequently, we recommend further research into this type of method.
We also observed that the Next-K approaches exhibit some interesting properties compared to autoregression. For instance, the entropy of label distributions in autoregression models decreases with increasing generation steps, as shown in Fig. 4.
@@
The likely reason for this behavior is the dependency of the model’s future output on its own past errors. Conversely, Next-K models demonstrate increasing entropy, which is expected as future events become harder to predict. This suggests that Next-K models have the potential to better predict the distributions of future labels, a factor that should be considered in future research.
@@ 6.4 The Maximum Sequence Length
We observed that long-horizon prediction quality largely depends on the maximum number of allowed predictions in autoregression and Next-K models. As shown in Fig. 5, the optimal number of predicted events is usually less then the maximum number of events in the horizon. For example, in Transactions dataset, the optimal length ranges from 3 to 5, depending on the method. This indicates that the quality of confidence estimation degrades for events further in the future, therefore it is beneficial to manually limit the number of predictions in the target metric.
@@
We therefore conclude that more attention should be paid to probability calibration between generation steps. This result may also indicate real progress in long-horizon prediction tasks, specifically the maximum horizon one can model accurately. T-mAP is the only metric capable of evaluating the optimal predicted sequence length, whereas the OTD metric does not depend on sequence length as it compares prefixes of fixed size.
7 Related Work
Temporal Point Processes. Traditional MTPP models make strong assumptions about the underlying generative process, resulting in Poisson and Hawkes processes [rizoiu2017hawkes-tutorial]. Over the last decade, there has been a shift towards achieving greater model flexibility by applying neural architectures. Some works have utilized classical RNNs [du2016rmtpp] and transformers [zuo2020thp], while others have proposed architectures with continuous time [mei2017nhp, rubanova2019ode]. In the HoTPP benchmark, we evaluate MTPP models with discrete and continuous time. Unlike other benchmarks, we also evaluate popular approaches from related fields, including next-event prediction models from event sequences [mcdermott2024eventstreamgpt, padhi2021tabgpt] and Next-K models from time series analysis [lim2021timeseriessurvey].
MTPP Evaluation. Previous benchmarks have primarily focused on medical data or traditional MTPP datasets. EventStream-GPT [mcdermott2024eventstreamgpt] and TemporAI [saveliev2023temporai] consider only medical data and do not implement methods from the MTPP field, despite their applicability. Early MTPP benchmarks, such as Tick [bacry2017tick] and PyHawkes [^1] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches and are unable to work with modern neural networks. PoPPy [xu2018poppy] and EasyTPP [xue2023easytpp] implement neural methods; however, these benchmarks cannot be extended to use Next-K approaches. PoPPy does not evaluate long-horizon prediction at all. This limitation is partially addressed in EasyTPP, which measures the OTD metric at the end of each sequence. In this work, we address the limitations of the OTD metric and introduce a superior alternative, T-mAP, along with an effective GPU-accelerated implementation.
8 Limitations
The training process for some methods is relatively slow. For example, a continuous time LSTM from the NHP method lacks an effective open-source GPU implementation, which could be developed in the future. Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training. Additionally, the list of implemented backbone architectures and losses can be extended, for example, by incorporating modern Transformer approaches[zuo2020thp] or Intensity-free training objective[shchur2019intensityfree].
9 Future Research Directions
In this work, we analyze long-horizon prediction evaluation and propose a new validation process. We strongly encourage future research to develop improved techniques for predicting future events and to establish simple baselines for better-measuring progress in the field. For example, Next-K models, which predict multiple future events without autoregression, show promising results in long-horizon prediction tasks. We suggest further exploration of these models. We also highlight the importance of label distribution estimation and emphasize the need to improve confidence estimation and calibration.
@@ 10 Conclusion
In this paper, we propose the HoTPP benchmark to assess the quality of multiple events forecasting. Our extensive evaluations using established datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction does not necessarily correlate with superior performance in horizon prediction. This finding underscores the need for benchmarks like HoTPP, which emphasize the importance of long-horizon accuracy and robustness. By shifting the focus from short-term to long-term predictive capabilities, HoTPP aims to drive the development of more sophisticated and reliable event sequence prediction models. This, in turn, has the potential to significantly enhance the practical applications of sequential event prediction in various domains, fostering innovation and improving decision-making processes across a wide range of industries.
@@
Appendix A Background on Marked Temporal Point Processes Modeling
[…]
The derivation is provided in [rizoiu2017hawkes-tutorial]. Different TPPs are characterized by their intensity function $\lambda(t_{i})$. In Poisson and non-homogeneous Poisson processes, the intensity function is independent of previous events, meaning event occurrences depend solely on external factors. In contrast, self-exciting processes are characterized by previous events increasing the intensity of future events. A notable example of a self-exciting process is the Hawkes process, in which each event linearly affects the future intensity:
@@ Appendix A Background on Marked Temporal Point Processes Modeling
where $\lambda_{0}(t_{i})\geq 0$ is the base intensity function, independent of previous events, and $\phi(x)\geq 0$ is the so-called memory kernel function. Given the intensity function $\lambda(t_{i})$, predictions are typically made by sampling or expectation estimation. Sampling is usually performed using the thinning algorithm, a specific rejection sampling approach. For details on the implementation of sampling, please refer to [rizoiu2017hawkes-tutorial].
@@
Recent research has focused on modeling complex intensity functions. Various neural network architectures have been adapted to address this problem. These approaches differ in the type of neural network used and the model of the intensity function. Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models, like NHP [mei2017nhp] and Neural ODEs [rubanova2019ode]. The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
@@
Intensity-free modeling. Some methods evaluate the next event time distribution without using intensity functions. For example, intensity-free [shchur2019intensityfree] represents the distribution as a mixture of Gaussians or through normalizing flows. Other approaches do not predict the distribution directly but instead, solve a regression problem using MAE or RMSE loss. However, it can be shown that both MAE and RMSE losses are closely related to distribution prediction. Specifically, RMSE is analogous to log-likelihood optimization with a Normal distribution, and MAE is similar to the log-likelihood of a Laplace distribution [bishop1994mdn]. In our experiments, we evaluated a model trained with MAE loss as an example of an intensity-free method.
@@
Rescoring with HYPRO. HYPRO [xue2022hypro] is an extension applicable to any sequence prediction method capable of sampling. HYPRO takes a pretrained generative model and trains an additional scoring module to select the best sequence from a sample. It is trained with a contrastive loss to distinguish between the generated sequence and the ground truth. HYPRO generates multiple sequences with a background model during inference and selects the best one by maximizing the estimated score. Although HYPRO is intended to improve quality compared to simple sampling, it is unclear whether HYPRO outperforms expectation-based prediction. In our work, we apply HYPRO to the outputs of the RMTPP model.
@@ Appendix B T-mAP Computation and Proofs
We have several correct and incorrect predictions for each threshold $h$ on the predicted label probability. These quantities define precision and recall values along with the total number of ground truth events. A prediction is correct if it has an assigned ground truth event within the required time interval $|t^{p}_{i}-t_{j}^{gt}|<\delta$. Note that each target can be assigned to at most one prediction. Therefore, we define a matching, i.e., the correspondence between predictions and ground truth events. The theorem states that the maximum size matching for each threshold $h$ can be constructed as a subset of an optimal matching between full sequences $\{\mathcal{S}_{p}^{i}\}_{i=1}^{n}$ and $\mathcal{S}^{i,l}_{gt}$, where optimal matching is a solution of the assignment problem with a bipartite graph, defined in Section 4.2.
@@ Appendix C HoTPP Benchmark Details
C.1 Methods
We implement popular methods from five groups: intensity-based, intensity-free, reranking, next-k, and statistical baselines.
Intensity-based methods. We implement the RMTPP[du2016rmtpp] and NHP[mei2017nhp] methods.
Intensity-free methods. As mentioned in Appendix A, this group includes intensity-free TPP and RMSE/MAE regression losses. We implement the MAE-CE approach, which uses MAE loss for time interval prediction and cross-entropy loss for labels.
Reranking approaches. We implement HYPRO[xue2022hypro] with the RMTPP backbone.
Next-K models. While not commonly used in the MTPP field, Next-K approaches are popular in time series modeling [lim2021timeseriessurvey]. These methods predict multiple future events at once, avoiding the need for autoregressive inference. The advantages of Next-K approaches include fast inference and stability, as predictions do not depend on potential errors from previous steps, unlike autoregression. The main limitation is a fixed prediction horizon, as the model cannot predict sequences of arbitrary length. However, applying a modified autoregressive approach can potentially address this limitation.
In our work, we implement a Next-K variant of the MAE-CE model, which is quite straightforward. In particular, given K predictions and K ground truth events, we just compute MAE and cross-entropy losses between the pairs of corresponding events from two sequences. A bit more complicated is the Next-K variant of the RMTPP approach, as it violates Hawkes’s assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1. So, RMTPP-K, unlike RMTPP, can’t model dependencies between the predicted K events. Despite this, RMTPP-K is still a strong model with top-performing results on the StackOverflow dataset.
In our work, we implement a Next-K variant of the MAE-CE model, which is straightforward. Given K predictions and K ground truth events, we compute MAE and cross-entropy losses between corresponding pairs of events from both sequences. The Next-K variant of the RMTPP approach is more complex, as it violates Hawkes assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1. Consequently, RMTPP-K, unlike RMTPP, cannot model dependencies between the predicted K events. Despite this, RMTPP-K still performs strongly, particularly on the StackOverflow dataset.
MostPopular baseline. This baseline analyzes the prefix sequence to identify the most frequent label and computes an average time step. The model generates constant output until the required duration (horizon) is reached. While this approach does not yield impressive results, it demonstrates moderate quality, especially in next-event prediction tasks.
Last-K baseline. This method extracts the required number of events from the end of the prefix and shifts timestamps to the current value. Despite its simplicity, this baseline demonstrates OTD quality comparable to neural approaches on small datasets.

### v2

@@ 6 Experiments
We conduct experiments on five datasets of varying sizes and origins: the Transactions dataset (AI-Academy, 2021), the MIMIC-IV medical dataset (Johnson et al., 2023), two social networks datasets, Retweet (Zhao et al., 2015) and StackOverflow (Jure, 2014), and the Amazon reviews dataset (Jianmo, 2018). Dataset statistics and evaluation parameters are provided in Appendix C.4.
@@
    Rule-based baselines. MostPopular Generates a constant output with the most popular label from the prefix and the average time step. The Last N baseline outputs N previous events with adjusted timestamps.
@@
    Intensity-free approaches. We implement IFTPP method, which combines mean absolute error (MAE) of the time step prediction with cross-entropy categorical loss for labels (Shchur et al., 2019; Padhi et al., 2021; McDermott et al., 2024).
@@
    Intensity-based approaches. We implement RMTPP (Du et al., 2016) as an example of the TPP approach with a traditional RNN. We add NHP (Mei & Eisner, 2017), based on a continuous time LSTM architecture. We also evaluate an AttNHP approach that utilizes a continuous time transformer model (Yang et al., 2022) and ODE (Rubanova et al., 2019).
@@
    Next-K approaches. We adapt IFTPP and RMTPP to predict multiple future events directly, implementing IFTPP-K and RMTPP-K, respectively. These approaches originate in time series analysis and have not been previously applied in the MTPP domain.
@@
    Reranking. We implement HYPRO (Xue et al., 2022), which generates multiple hypotheses and selects the best sequence using a contrastive approach.
Additional details on these methods are provided in Appendix A, with training specifics outlined in Appendix D. Hyperparameters for both the methods and metrics are presented in Appendix E. The main evaluation results are shown in Table 1. In the following sections, we will discuss these results from various perspectives and offer additional analysis of the methods’ behavior.
6.1 Metrics Comparison
The evaluation results in Table 1 show that the OTD metric can yield low values even for simple rule-based baselines. For instance, the Last baseline achieves the lowest OTD values on the Retweet and Amazon datasets. In contrast, the T-mAP metric more clearly distinguishes between rule-based baselines and deep learning models. This difference arises from the ability to measure model confidence. OTD evaluates predicted labels alone, while T-mAP measures average precision by adjusting logits thresholds. Unlike rule-based baselines, deep models provide confidence scores, whereas baselines output only labels, resulting in low T-mAP scores for simple baselines.
The HYPRO and NHP methods maximize T-mAP, while the lowest OTD values are achieved by statistical baselines, IFTPP, and RMTPP. The difference between the two metrics lies in balancing the importance of event order versus time prediction quality. OTD evaluates the first 5-10 events, and to minimize OTD, a model must correctly identify the events that constitute the beginning of the sequence. IFTPP and RMTPP jointly predict the next event’s time and label, leading to better ordering. In contrast, methods like NHP, AttNHP, and ODE predict time steps independently for each event type. When the differences between timestamps are small, these methods can result in random ordering, negatively affecting the OTD score. T-mAP, on the other hand, is less sensitive to event order as long as the time predictions are accurate.
The HYPRO method, designed to differentiate between ground truth and predicted sequences using a discriminator model, achieves high T-mAP scores. This highlights T-mAP’s ability to reward models that generate realistic sequences. We also observed that methods with high T-mAP scores often have low next-event prediction accuracy. For instance, on the Retweet, Amazon, and MIMIC-IV datasets, high long-horizon performance is often accompanied by low next-event prediction scores. This is particularly true for the HYPRO method, whose loss function is heavily focused on long-horizon quality. The contrast on the MIMIC-IV dataset can be attributed to the large number of events with identical timestamps, making precise event ordering more difficult and impacting both next-event and OTD scores.
@@ 6.2 Intensity-Based vs Intensity-Free
Previous works mostly compared either intensity-based neural methods (Du et al., 2016; Xue et al., 2022) or intensity-free approaches (Padhi et al., 2021; McDermott et al., 2024). The only exception is IFTPP (Shchur et al., 2019), which did not compare with NHP and included only test set likelihood as an evaluation metric. This raises the question: which approach is superior? We compare the intensity-free IFTPP method with intensity-based RMTPP, NHP, and ODE approaches. Table 1 shows that the intensity-free IFTPP method excels in the next-event MAE prediction, as it optimizes this metric. In other scenarios, the results vary between datasets. Therefore, we conclude that there is no preferred solution, and both approaches warrant attention in future research.
@@ 6.3 Autoregression vs Direct Horizon Prediction
We compared the IFTPP and RMTPP methods with their Next-K counterparts. Our results indicate that autoregressive models generally perform better regarding next-event prediction quality, while Next-K approaches improve T-mAP scores on all datasets except Transactions. The OTD-based rankings vary across datasets, with no single approach consistently outperforming the others. Despite these findings, little effort has been made to adapt Next-K models to MTPPs, suggesting future research in this area.
We also observed that the Next-K approaches exhibit some interesting properties compared to autoregression. For instance, the entropy of label distributions in autoregression models decreases with increasing generation steps, as shown in Figure 5. The likely reason for this behavior is the dependency of the model’s future output on its past errors. Conversely, Next-K models demonstrate increasing entropy, which is expected as future events become harder to predict. This suggests that Next-K models have the potential to predict better the distributions of future labels, a factor that should be considered in future research.
@@ 6.4 The Maximum Sequence Length
We observed that long-horizon prediction quality largely depends on the maximum number of allowed predictions in autoregression and Next-K models. As shown in Fig. 6, the optimal number of predicted events is usually less than the maximum number of events on the horizon. For example, in the Transactions dataset, the optimal length ranges from 3 to 5, depending on the method. This indicates that the quality of confidence estimation degrades for events further in the future, therefore it is beneficial to limit the number of predictions manually.
@@
Therefore, we conclude that greater attention should be given to probability calibration between generation steps. This finding also highlights progress in long-horizon prediction tasks, particularly determining the maximum horizon that can be accurately modeled. T-mAP is the only metric capable of evaluating the optimal predicted sequence length, while the OTD metric remains unaffected by sequence length as it compares fixed-size prefixes.
7 Limitations and Future Work
The proposed benchmark has several shortcomings. The training process for some methods is relatively slow. For example, a continuous time LSTM from the NHP method lacks an effective open-source GPU implementation, which could be developed in the future. Autoregression can also be optimized using specialized GPU implementations, which would significantly impact HYPRO training. Additionally, the list of implemented backbone architectures and losses can be extended, for example, by incorporating hybrid (Deshpande et al., 2021) and diffusion models (Zhou et al., 2023).
Our benchmark encourages future research to develop improved techniques for predicting future events and establish simple baselines for better measuring progress in the field. For example, Next-K models, which predict multiple future events without autoregression, show promising results in long-horizon prediction tasks. We suggest further exploration of these models. We also highlight the importance of label distribution estimation and emphasize the need to improve confidence estimation and calibration.
The T-mAP metric can be applied in domains beyond event sequences, like action recognition (Su & Hua, 2017). T-mAP can potentially offer better timestamp evaluation in these domains than evaluating indices with OPW. T-mAP can also provide a more natural hyperparameter selection regarding the modeling horizon and maximum allowed time error. Our theoretical justification of T-mAP can potentially be adapted to the mAP estimation algorithms in computer vision.
@@ 8 Conclusion
In this paper, we propose the HoTPP benchmark to assess the quality of long-horizon events forecasting. Our extensive evaluations using established datasets and various predictive models reveal a critical insight: high accuracy in next-event prediction does not necessarily correlate with superior performance in horizon prediction. This finding underscores the need for benchmarks like HoTPP, which emphasize the importance of long-horizon accuracy and robustness. By shifting the focus from short-term to long-term predictive capabilities, HoTPP aims to drive the development of more sophisticated and reliable event sequence prediction models. This, in turn, has the potential to significantly enhance the practical applications of sequential event prediction in various domains, fostering innovation and improving decision-making processes across a wide range of industries.
Appendix A Background on Marked Temporal Point Processes Modeling
Intensity-based approaches. Modeling the probability density function (PDF) $f^{*}(t_{i})=f(t_{i}|t_{1},\dots,t_{i-1})$ for the next event time is typically challenging, as it requires the additional constraint that its integral equals one. Instead, the non-negative intensity function $\lambda(t_{i})\geq 0$ is usually modeled. The following equation gives the relationship between the PDF and the intensity function:
@@ Appendix A Background on Marked Temporal Point Processes Modeling
The derivation is provided in (Rizoiu et al., 2017). Different TPPs are characterized by their intensity function $\lambda(t_{i})$. In Poisson and non-homogeneous Poisson processes, the intensity function is independent of previous events, meaning event occurrences depend solely on external factors. In contrast, self-exciting processes are characterized by previous events increasing the intensity of future events. A notable example of a self-exciting process is the Hawkes process, in which each event linearly affects the future intensity:
@@
where $\lambda_{0}(t_{i})\geq 0$ is the base intensity function, independent of previous events, and $\phi(x)\geq 0$ is the so-called memory kernel function. Given the intensity function $\lambda(t_{i})$, predictions are typically made by sampling or expectation estimation. Sampling is usually performed using the thinning algorithm, a specific rejection sampling approach. For details on the implementation of sampling, please refer to (Rizoiu et al., 2017).
@@
Recent research has focused on modeling complex intensity functions. Various neural network architectures have been adapted to address this problem. These approaches differ in the type of neural network used and the model of the intensity function. Neural architectures range from simple RNNs and Transformers to specially designed continuous-time models like NHP (Mei & Eisner, 2017) and Neural ODEs (Rubanova et al., 2019). The intensity function between events can be modeled as a sum of intensities from previous events, as in Hawkes processes, or by directly predicting the inter-event intensity given the context embedding.
@@
Intensity-free modeling. Some methods evaluate the next event time distribution without using intensity functions. For example, intensity-free (Shchur et al., 2019) represents the distribution as a mixture of Gaussians or through normalizing flows. Instead of predicting the distribution directly, other approaches solve a regression problem using MAE or RMSE loss. However, it can be shown that both MAE and RMSE losses are closely related to distribution prediction. Specifically, RMSE is analogous to log-likelihood optimization with a Normal distribution, and MAE is similar to the log-likelihood of a Laplace distribution (Bishop, 1994). In our experiments, we evaluated a model trained with MAE loss as an example of an intensity-free method.
@@
Rescoring with HYPRO. HYPRO (Xue et al., 2022) is an extension applicable to any sequence prediction method capable of sampling. HYPRO takes a pretrained generative model and trains an additional scoring module to select the best sequence from a sample. It is trained with a contrastive loss to distinguish between the generated sequence and the ground truth. HYPRO generates multiple sequences with a background model during inference and selects the best one by maximizing the estimated score. Although HYPRO is intended to improve quality compared to simple sampling, it is unclear whether HYPRO outperforms expectation-based prediction. In our work, we apply HYPRO to the outputs of the RMTPP model.
Next-K models. While not commonly used in the MTPP field, Next-K approaches are popular in time series modeling (Lim & Zohren, 2021). These methods predict multiple future events at once, avoiding the need for autoregressive inference. The advantages of Next-K approaches include fast inference and stability, as predictions do not depend on potential errors from previous steps, unlike autoregression. The main limitation is a fixed prediction horizon, as the model cannot predict sequences of arbitrary length. However, applying a modified autoregressive approach can potentially address this limitation.
In our work, we implement a Next-K variant of the IFTPP model, which is straightforward. Given K predictions and K ground truth events, we compute MAE and cross-entropy losses between corresponding pairs of events from both sequences. The Next-K variant of the RMTPP approach is more complex, as it violates Hawkes assumptions: the $i$-th prediction doesn’t depend on predictions $11\dots i-$ 1. Consequently, RMTPP-K, unlike RMTPP, cannot model dependencies between the predicted K events. Despite this, RMTPP-K still performs strongly, particularly on the StackOverflow dataset.
@@ Appendix B T-mAP Computation and Proofs
We have several correct and incorrect predictions for each threshold $h$ on the predicted label probability. These quantities define precision and recall values along with the total number of ground truth events. A prediction is correct if it has an assigned ground truth event within the required time interval $|t^{p}_{i}-t_{j}^{gt}|\leq\delta$. Note that each target can be assigned to at most one prediction. Therefore, we define a matching, i.e., the correspondence between predictions and ground truth events. The theorem states that the maximum size matching for each threshold $h$ can be constructed as a subset of an optimal matching between full sequences $\{\mathcal{S}_{p}^{i}\}_{i=1}^{n}$ and $\mathcal{S}^{i,l}_{gt}$, where optimal matching is a solution of the assignment problem with a bipartite graph, defined in Section 4.2.
@@
Calibration dependency. Calibration influences the weights assigned to the edges of the graph $\mathcal{G}$. T-mAP computation involves two key steps for each class label: matching and AP estimation. While average precision (AP) is invariant to monotonic transformations of predicted class logits, the matching step is only invariant to linear transformations. Specifically, the optimal matching seeks to minimize the total weight in the following form:
$C=\operatorname*{arg\,min}\limits_{m\in M(\mathcal{G})}\sum\limits_{(i,j)\in m}(-s^{p}_{i}).$ (9)
A linear transformation of logits with a positive scaling factor will adjust the total weight accordingly, but the optimal matching will remain unchanged. Since the matching is performed independently for each class, we conclude that T-mAP is invariant to linear calibration.
@@ Appendix C HoTPP Benchmark Details
C.1 Architecture
HoTPP incorporates best practices from extensible and reproducible ML pipelines. It leverages PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) as the core training library, ensuring reproducibility and portability across various computing environments. Additionally, HoTPP utilizes the Hydra configuration library (Yadan, 2019) to enhance extensibility. The overall architecture is illustrated in Figure 8. HoTPP supports both discrete-time and continuous-time models, as well as RNN and Transformer architectures. Implementing a new method requires only the addition of essential components, while the rest can be configured through Hydra and configuration files.
Figure 8: The architecture of the HoTPP library.